OpenStudy (anonymous):

Which equation is a quadratic equation? 5x + y = -2(x + 5) y - 9 = (x2 + 1)(5x - 1) + 3 y + 7x = -3x(7 - 2x) + 11 3x3 - y = (3x - 2)(x + 5)

OpenStudy (anonymous):

help please :(

jimthompson5910 (jim_thompson5910):

you want something that has an x^2 term

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

and you want this to be the largest term (ie you don't want something with x^3, x^4, etc)

jimthompson5910 (jim_thompson5910):

so choice D is out because 3x^3 tells us that we're dealing with a cubic (not a quadratic)

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

what do you get when you expand out choice B?

OpenStudy (anonymous):

uhhhh

jimthompson5910 (jim_thompson5910):

y - 9 = (x^2 + 1)(5x - 1) + 3 y - 9 = x^2(5x - 1) + 1(5x-1) + 3 y - 9 = 5x^3 - x^2 + 5x-1 + 3 so what type of equation is this?